OpenStudy (anonymous):

Can someone help me please? What is the perimeter of a triangle with the given side lengths? s1 = x + 4 cm s2 = 3x + 1 cm s3 = 7x + 3 cm

OpenStudy (alexandervonhumboldt2):

to find the perimeter just add sides: s1+s2+s3=x+4+3x+1+7x+3=combine like terms and simplify

OpenStudy (anonymous):

oh! okay, is that it?

OpenStudy (alexandervonhumboldt2):

well you can combine x, 7x and 3x and then combine 4, 1 and 3 what is x+7x+3x?

OpenStudy (anonymous):

10x?

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

nearly

OpenStudy (anonymous):

okay, can you tell me what i am doing wrong?

OpenStudy (alexandervonhumboldt2):

well, why you left x? 7x+3x+x=x*(7+3+1)=11x

OpenStudy (alexandervonhumboldt2):

ok so we have 11x now

OpenStudy (anonymous):

ok

OpenStudy (alexandervonhumboldt2):

lets combine 4, 1 and 3 now what is 4+1+3=?

OpenStudy (anonymous):

8

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

so the pertimeter is 11x+8

OpenStudy (anonymous):

oh! i get it now!

OpenStudy (alexandervonhumboldt2):

:)

OpenStudy (anonymous):

thanks!